OpenStudy (anonymous):

-7x – 4y = -15 -7x – 3y = -13

OpenStudy (anonymous):

Do you need to find x and y?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Solve one of the equation for x or y and sub it into the other one to replace the solved for variable. So, if you were to solve the 1st equation for y, it would be y= (15/4)-(7/4)x

OpenStudy (anonymous):

the answerwould be 3

OpenStudy (anonymous):

Yup.

OpenStudy (anonymous):

what about the 2 equation

OpenStudy (anonymous):

to find y make x =0 to find x plug the number that you found for y in the equation ..

OpenStudy (anonymous):

i got 1

OpenStudy (anonymous):

for the first eq it would be -7(0)-4y=-15 then -4=-15 , so y = -15 /-4 which gives 15/4

OpenStudy (anonymous):

since you have y = 15/4 plug this in the eq to find x

OpenStudy (anonymous):

so what are the two ansers

OpenStudy (anonymous):

15/4 7

OpenStudy (anonymous):

can i see the question ?

OpenStudy (anonymous):

solve a system using elimination and those two equations

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

first step you should cancel or remove y or x in both equations

OpenStudy (anonymous):

for example you have :-7x – 4y = -15 -7x – 3y = -13 .. since x in both equations have the same number its easier that y to cancel .. right ?

OpenStudy (anonymous):

but before you cancel x you have to multiply one of the equations by +1 to get one x has a positive number and the other one has a negative number which is -7 and +7

OpenStudy (anonymous):

then

OpenStudy (anonymous):

its the same as the following example 2x + y = 9 3x – y = 16 Note that, if I add down, the y's will cancel out. So I'll draw an "equals" bar under the system, and add down: 2x + y = 9 3x – y = 16 5x = 25 Now I can divide through to solve for x = 5, and then back-solve, using either of the original equations, to find the value of y. The first equation has smaller numbers, so I'll back-solve in that one: 2(5) + y = 9 10 + y = 9 y = –1 Then the solution is (x, y) = (5, –1).

OpenStudy (anonymous):

the answer was 1,2